OpenStudy (anonymous):

A triangle has vertices P(1, 3), Q(3, 5), and R(6, 2). The slope of the median to QR is?

OpenStudy (anonymous):

@acxbox22

OpenStudy (acxbox22):

plot the three coordinates

OpenStudy (anonymous):

don't you just find the slope of QR since slope doesn't change in a line

OpenStudy (jdoe0001):

lemme fix that

OpenStudy (acxbox22):

you find the midpoint of QR then you find the slope of the point opposite side QR and the midpoint of QR

OpenStudy (anonymous):

ok i graphed it

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) Q&({\color{red}{ 3}}\quad ,&{\color{blue}{ 5}})\quad % (c,d) R&({\color{red}{ 6}}\quad ,&{\color{blue}{ 2}}) \end{array}\qquad % coordinates of midpoint \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\) anyhow, find the median point and use that with the point P to get the slope for that

OpenStudy (anonymous):

why would i use p

OpenStudy (jdoe0001):



OpenStudy (anonymous):

yep except r should be lower than p

OpenStudy (jdoe0001):

hmmm "not drawn to scale" :P

OpenStudy (anonymous):

wait i understand what the question is asking now

OpenStudy (jdoe0001):

but you're correct, it is lower than P

OpenStudy (jdoe0001):

the slope of midpoint of QR and P

OpenStudy (anonymous):

midpoint is (-3/2,3/2)

OpenStudy (anonymous):

1.5/2.5

OpenStudy (anonymous):

but thats not an answer

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

I guessed and somehow the answer is 1/7?

OpenStudy (jdoe0001):

well.. your midpoint is .... off for one I got something else

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

wait nevermind I screwed up my numbers oops

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) Q&({\color{red}{ 3}}\quad ,&{\color{blue}{ 5}})\quad % (c,d) R&({\color{red}{ 6}}\quad ,&{\color{blue}{ 2}}) \end{array}\qquad % coordinates of midpoint \left(\cfrac{{\color{red}{ 6}} + {\color{red}{ 3}}}{2}\quad ,\quad \cfrac{{\color{blue}{ 2}} + {\color{blue}{ 5}}}{2} \right)\implies ?\)

OpenStudy (anonymous):

Thank you I did it again and got 1/7

OpenStudy (anonymous):

yeah thats what i did the second time

OpenStudy (anonymous):

(1.5,2.5)

OpenStudy (jdoe0001):

well... yes but it happens to give a slope of 1/7, yes but with the correct midpoint

OpenStudy (anonymous):

Yep I messed up, but i got to it, thank you for helping me

OpenStudy (jdoe0001):

yw